# Jacquet modules of Tate cohomology and base change lifting

### v1

@@ Abstract.
Let $G$ be a connected reductive algebraic group defined over a non-Archimedean local field $F$ of characteristic $p$. Let $\ell$ be a prime number with $\ell\neq p$. Let $\Gamma$ be a group of automorphisms of $G$ of order $\ell$. Let $\Pi$ be an $\ell$-modular irreducible smooth representation of $G(F)$ which extends as a representation of $G(F)\rtimes\Gamma$. We show that the Jacquet modules of the Tate cohomology groups $\widehat{H}^{i}(\Gamma,\Pi)$ are sub modules of certain Jacquet modules of $\Pi$. When $\Gamma$ is induced by a Galois automorphism, we use this result on Jacquet modules to show that $\widehat{H}^{i}(\Gamma,\Pi)$ is a finite length representation. We further use the Jacquet-modules to study some irreducibility of Tate cohomology in the cyclic base change lifting of cuspidal representations. The novelty in this article is that we do not assume that the base change lifting is cuspidal. Later in this article, we study these results for the finite fields case where we obtain a much precise description of the Tate cohomology groups. Finally, we end this article by studying the Tate cohomology of the $-\ell$ reduction of the unipotent cuspidal representation of ${\rm Sp}_{4}(\mathbb{F}_{q^{\ell}})$ for the action of ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$.
@@ 1. Introduction
Let $G$ be a connected reductive group defined over a $p$-adic field $F$. Let $\Gamma$ be a group of automorphisms of $G$ defined over $F$, and of cardinality a prime number $\ell$, and we assume that $\ell\neq p$ in this article. We fix a generator $\sigma$ of $\Gamma$, and we denote by $G^{\sigma}$ the connected component of the fixed point subgroup of $G$. Let $(\Pi,V)$ be an $\ell$-modular smooth irreducible representation of $G(F)$ such that $\Pi$ extends to a representation of $G(F)\rtimes\Gamma$. Treumann and Venkatesh, in their work on $-\ell$ functoriality ([TV16]), proposed some remarkable conjectures for the representation theory of $p$-adic groups. One of their conjectures predict that the Tate cohomology groups $\widehat{H}^{i}(\Gamma,\Pi)$ are finite length representations of $G^{\sigma}(F)$, for $i\in\{0,1\}$, and they also proposed that the sub-quotients of $\widehat{H}^{i}(\Gamma,\Pi)$ must be related to $\Pi$ via $\ell$-modular functoriality in the sense of Langlands, and they called it the linkage principle. In our previous work [ND24], we proved that $\widehat{H}^{i}(\Gamma,\Pi)$ is a finite length representation for the case where $G={\rm Res}_{E/F}{\rm GL}_{n}$ and $\sigma$ is the automorphism of $G$ induced by a generator of the cyclic group ${\rm Gal}(E/F)$ of order $\ell$, and we showed the linkage principle holds when $\ell$ does not divide the pro-order of ${\rm GL}_{n-1}(F)$.
@@
In this article, we prove the finiteness part of Treumann–Venkatesh conjecture for the case where $\Gamma$ is induced by Galois automorphisms, i.e., the case where $G$ is ${\rm Res}_{E/F}H$ where $H$ is a connected reductive group defined over $F$, the extension $E/F$ is a cyclic Galois extension of degree $\ell$, and $\Gamma$ is induced by the Galois group ${\rm Gal}(E/F)$. The main technique is to control of Jacquet modules and twisted Jacquet modules of the Tate cohomology groups. We also use our results on Jacquet modules to further show that the $\Gamma$-Tate cohomology of a ${\rm GL}_{n}(E)\rtimes\Gamma$-stable lattice in the cyclic base change lift to ${\rm GL}_{n}(E)$, of a smooth $\ell$-adic cuspidal representation of ${\rm GL}_{n}(F)$ is irreducible and cuspidal. In the depth-zero case, using local methods, we obtain the precise description of Tate cohomology of a base change lift. We also obtain some results on the Tate cohomology of invariant lattices in Shintani lifts for ${\rm GL}_{n}$ over finite fields.
@@
In two papers [Fen24] and [Fen23], Feng showed that, under mild restrictive hypothesis on the groups involved, the Genestier–Lafforgue correspondence and the Fargues–Scholze correspondence is compatible with the linkage principle. However, the explicit computation of Tate cohomology, in particular, admissibility, the non-vanishing of Tate cohomology $\widehat{H}^{i}(\Gamma,\Pi)$, are not well understood. And we attempt to answer some of these questions in this article.
@@
We will now precisely describe our principal results. The first result of this paper is to show the Tate cohomology groups are finitely generated representations. We prove that
Theorem 1.1.
@@
Let $G$ be a connected reductive algebraic group over a non-Archimedean local field $F$. Let $\Gamma\subset{\rm Aut}(G)(F)$ be a group of automorphisms of order $\ell$. Let $(\Pi,V)$ be a finite length representation of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$. Then the Tate cohomology $\widehat{H}^{i}(\Gamma,\Pi)$ is a finitely generated representation of $G^{\sigma}(F)$.
The proof the above theorem uses the $\ell$-modular version of Schneider and Stuhler resolutions ([SS97, Theorem II.3.1]) of smooth irreducible representations proved by Vignéras ([Vig97, Proposition 2.6]). In what follows, we use the hypothesis and notations of the above theorem. Now, consider a parabolic subgroup $P$ of $G$ which is $\sigma$ stable such that $P^{\sigma}$ is a parabolic subgroup of $G^{\sigma}$. Let $U$ be the unipotent radical of $P$ and let $N$ be the unipotent radical $P^{\sigma}$. Let $\Psi$ be a character of $U(F)$. We will show that
$\widehat{H}^{i}(\Gamma,\Pi)_{N(F),\Psi}\hookrightarrow\widehat{H}^{i}(\Gamma,\Pi_{U(F),\Psi}).$ (1.1)
@@
This article, in essence, is about various applications of the above injective map. First of which is the following theorem.
Theorem 1.2.
@@
Let $G$ be a connected reductive group over a non-Archimedean local field $F$, and let $E/F$ be a cyclic Galois extension of degree $\ell$. Let $\Pi$ be a finite length representation of $\overline{\mathbb{F}}_{\ell}[G(E)\rtimes{\rm Gal}(E/F)]$. Then the Tate cohomology $\widehat{H}^{i}({\rm Gal}(E/F),\Pi)$ is a finite length representation of $\overline{\mathbb{F}}_{\ell}[G(F)]$.
@@
We prove the above theorem using induction on the dimension of $G$. In the induction argument, we need to control the action of the centre of $G^{\sigma}$ along with the exponents in Jacquet modules with respect to parabolic subgroups of $G^{\sigma}$. For this reason, in this article, we restrict to the case where $\sigma$ is induced by a Galois automorphism. In the Galois automorphism case, for a Levi subgroup $M$ defined over $F$, the action of the centre of $M(F)$ on the ${\rm Gal}(E/F)$- Tate cohomology of the Jacquet modules of $\Pi$ is via the restriction of the centre of $M(E)$. In other words, we are using the fact that $Z(M(E))\cap G(F)=Z(M(F))$. In the general case, i.e., when $\sigma$ is not necessarily induced by a Galois automorphism, we note that $Z(G^{\sigma})$ can be quite large compared to $G^{\sigma}$. For instance, $\sigma$ could be an inner automorphism induced by a regular element of order $\ell$, and in this case, $G^{\sigma}$ is a torus.
@@
We will now assume that $G$ is ${\rm Res}_{E/F}{\rm GL}_{n}$, where $\sigma$ is an automorphism of $G$, induced by a generator of ${\rm Gal}(E/F)$. Here, $E/F$ is a cyclic Galois extension of degree $\ell$. Let $\mathcal{K}$ be the maximal unramified extension of $\mathbb{Q}_{\ell}$ and $\Lambda$ be the ring of integers of $\mathcal{K}$. Let $\Pi$ be an absolutely irreducible, integral, representation of ${\rm GL}_{n}(E)$, defined over $\mathcal{K}$ such that $\Pi^{\sigma}\simeq\Pi$. Let $\pi$ be an integral $\ell$-adic irreducible cuspidal representation of ${\rm GL}_{n}(F)$ such that $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lifting of $\pi$. Note that $\Pi$ is not necessarily a cuspidal representation. We show the following result on any $\Lambda[{\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)]$ stable lattice in $\Pi$.
Theorem 1.3.
@@
With the Hypothesis on $\Pi$ and $\pi$ as above, let $\mathcal{L}$ be a $\Lambda[{\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)]$ stable lattice in $\Pi$. Then, the space $\widehat{H}^{0}({\rm Gal}(E/F),\mathcal{L})$ is an irreducible cuspidal representation of ${\rm GL}_{n}(F)$ and $\widehat{H}^{1}({\rm Gal}(E/F),\mathcal{L})$ is trivial.
@@
For proving the above result, we first prove that the Tate cohomology group of a $\Lambda[{\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)]$ stable lattice in $\Pi$ has a unique generic sub-quotient. Infact this holds true in greater generality. Let $G$ be a split reductive group defined over $F$. Let $\Pi$ be a smooth irreducible $\ell$-modular representation of $G(E)$ such that $\Pi$ is stable under the action of ${\rm Gal}(E/F)$. Let $B$ be a Borel subgroup of $G$ defined over $F$, and let $U$ be the unipotent radical of $B$. Let $\psi:U(E)\rightarrow\overline{\mathbb{F}}_{\ell}^{\times}$ be a non-degenerate, ${\rm Gal}(E/F)$ invariant character such that $\psi$ restricts to a non-degenerate character on $U(F)$. If $\Pi$ is $(U(E),\psi)$-generic, then we show that $\widehat{H}^{i}({\rm Gal}(E/F),\Pi)$ has a unique $(U(F),\psi)$-generic sub-quotient (Theorem 3.5). To finally prove the theorem, we use the injection (1.1) and prove the right-hand side of the equation (1.1) is trivial by using the multiplicity free property of some Jacquet modules in the case where the base change lift is not cuspidal.
@@
We then consider the case of finite fields. The analogue of the above theorem holds true in the case where $\Pi$, a representation of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$, is a Shintani lift of a cuspidal representation of ${\rm GL}_{n}(\mathbb{F}_{q})$, and we use the character formula to show that $\widehat{H}^{0}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})$ is isomorphic to the Frobenius twist of the $-\ell$ reduction of $\pi$ (Theorem 4.4). The first Tate cohomology group in this case is proved to be trivial.
@@
We end this article by considering the example of $\theta_{10}$, the unipotent cuspidal irreducible representation of the symplectic group ${\rm Sp}_{4}(\mathbb{F}_{q})$. To be precise, let $\pi_{n}$ be the unique cuspidal unipotent representation of ${\rm Sp}_{4}(\mathbb{F}_{q^{n}})$. Note that the $-\ell$ reduction of $\pi_{n}$ is irreducible and is denoted by $r_{\ell}(\pi_{n})$. For a vector space $W$ over the field $\overline{\mathbb{F}}_{\ell}$, we denote by $W^{(\ell)}$ the $\overline{\mathbb{F}}_{\ell}$-vector space where the scalar action is twisted by the Frobenius automorphism of the field $\overline{\mathbb{F}}_{\ell}$. We will show the following theorem.
Theorem 1.4.
@@
The above theorem seems to suggest that for all classical groups $G$ admitting a cuspidal unipotent representation, the zeroth Tate cohomology for the action of ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$ on the $-\ell$ reduction of a cuspidal unipotent representation of $G(\mathbb{F}_{q^{\ell}})$ is isomorphic to the Frobenius twist of the $-\ell$ reduction of the cuspidal unipotent representation of $G(\mathbb{F}_{q})$.
@@
    We use the standard notations attached to any non-Archimedean local field $F$: the ring of integers of $F$ is $\mathfrak{o}_{F}$, the maximal ideal of $\mathfrak{o}_{F}$ is $\mathfrak{p}_{F}$, and the residue field $\mathfrak{o}_{F}/\mathfrak{p}_{F}$ is $k_{F}$. The field $k_{F}$ has $q_{F}$ elements and $q_{F}$ is a power of a prime $p$. Let $G$ be a connected reductive algebraic group defined over $F$.
@@
    Let $\sigma\in{\rm Aut}(G)$ be an element of prime order $\ell$ and $\ell\neq p$. Let $\Gamma$ be the group generated by $\sigma$. Let $G^{\sigma}$ be the connected component of the fixed points of $\sigma$, and it is assumed to be reductive.
@@
    A representation $(\pi,V)$ of $G(F)$ is said to be smooth if every vector $v\in V$ is fixed by a compact open subgroup of $G(F)$. Let $P$ be an $F$-rational parabolic subgroup of $G$, the unipotent radical is denoted by $U$, and let $M$ be its Levi subgroup defined over $F$. For a smooth representation $(\tau,W)$ of $M(F)$, we denote by $i_{P(F)}^{G(F)}(\tau)$ the normalised parabolic induction of $\tau$ to $G(F)$. We use the notation $r_{U(F)}(\pi)$ for the normalised Jacquet module of a smooth representation $(\pi,V)$ of $G(F)$.
@@
    In this article, a smooth representation $(\pi,V)$ of $G(F)$ is cuspidal if $r_{U(F)}(\pi)=0$, where $U$ is the unipotent radical of any proper $F$-rational parabolic subgroup $P$ of $G$ For a closed subgroup $H_{1}$ of a reductive $p$-adic group $H$, we will denote by ${\rm Ind}_{H_{1}}^{H}$ and ${\rm ind}_{H_{1}}^{H}$ the smooth induction functor and the compact induction functor respectively.

### v2

@@ Abstract.
Let $G$ be a connected reductive group defined over a non-Archimedean local field $F$ of residue characteristic $p$. Let $\ell$ be a prime number distinct from $p$. Let $E$ be a cyclic Galois extension of $F$ with $[E:F]=\ell$. Let $\Pi$ be a finite length $\overline{\mathbb{F}}_{\ell}$-representation (or an $\ell$-modular representation) of $G(E)\rtimes{\rm Gal}(E/F)$. In this context, we prove a conjecture of Treumann and Venkatesh which predicts that the Tate cohomology groups $\widehat{H}^{i}({\rm Gal}(E/F),\Pi)$ are finite length representations of $G(F)$. We discuss the explicit computation of these Tate cohomology groups when $G$ is ${\rm GL}_{n}$ and $\Pi$ is obtained as a base change lifting of a depth-zero cuspidal representation of ${\rm GL}_{n}(F)$. The primary novelty from our previous work is that we treat the case where $\Pi$ is possibly non-cuspidal. We also study the ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$-Tate cohomology groups of the $-\ell$ reduction of the unipotent cuspidal representation of ${\rm Sp}_{4}(\mathbb{F}_{q^{\ell}})$.
@@ 1. Introduction
Let $G$ be a connected reductive algebraic group defined over a $p$-adic field $F$. Let $\ell$ be a prime number distinct from $p$. Let $\sigma$ be an order $\ell$ automorphism of $G$, defined over $F$, and let $\Gamma$ be the cyclic group generated by $\sigma$. We denote by $G^{\sigma}$ the connected component of the subgroup of $\sigma$-fixed points of $G$. Let $\Pi$ be an $\ell$-modular smooth irreducible representation of $G(F)$ such that $\Pi$ extends to a representation of $G(F)\rtimes\Gamma$. Treumann and Venkatesh, in their work on $-\ell$ functoriality ([TV16, Conjectures 6.3]), proposed some remarkable conjectures relating the $\ell$-modular irreducible smooth representations of the $p$-adic groups $G(F)$ and $G^{\sigma}(F)$. Their conjectures predict that the Tate cohomology groups $\widehat{H}^{i}(\Gamma,\Pi)$ are finite length representations of $G^{\sigma}(F)$, for $i\in\{0,1\}$, and they also proposed that the sub-quotients of $\widehat{H}^{i}(\Gamma,\Pi)$ must be related to $\Pi$ via $\ell$-modular functoriality in the sense of Langlands, and they called it the linkage principle. In our previous work [ND24], we proved that $\widehat{H}^{i}(\Gamma,\Pi)$ is a finite length representation for the case where $G={\rm Res}_{E/F}{\rm GL}_{n}$ and $\sigma$ is the automorphism of $G$ induced by a generator of the cyclic group ${\rm Gal}(E/F)$ of order $\ell$, and we showed that the linkage principle holds when $\ell$ does not divide the pro-order of ${\rm GL}_{n-1}(F)$.
@@
In this article, we first prove the finiteness part of Treumann–Venkatesh conjecture for the case where $\sigma$ is induced by a Galois automorphism, i.e., the case where $G$ is ${\rm Res}_{E/F}H$ and $H$ is a connected reductive algebraic group defined over $F$, the extension $E/F$ is a cyclic Galois extension of degree $\ell$, and $\sigma$ is induced by a generator of the Galois group ${\rm Gal}(E/F)$. The main technique is to control the Jacquet modules and twisted Jacquet modules of the Tate cohomology groups. We also use the fact that subrepresentations of finitely generated $\ell$-modular representations of a $p$-adic group are finitely generated; this is a deep result proved recently in [DHKM24]. Let $\Pi$ be the base change lift to ${\rm GL}_{n}(E)$ of an integral cuspidal irreducible representation $\pi$ of ${\rm GL}_{n}(F)$ defined over the maximal unramified extension $\mathcal{K}$ of $\mathbb{Q}_{\ell}$. We further show that $\widehat{H}^{0}({\rm Gal}(E/F),\mathcal{L})$ is irreducible and a cuspidal representation, for any ${\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)$-stable $\mathbb{Z}_{\ell}^{\rm un}$-lattice $\mathcal{L}$ in $\Pi$. In the depth-zero case, using local methods, we obtain the precise description of Tate cohomology of a base change lift. These results on depth-zero representations are obtained by studying the Tate cohomology of invariant lattices in Shintani lifts for ${\rm GL}_{n}$ over finite fields.
@@
In two papers [Fen23] and [Fen24], Feng showed that, under mild restrictive hypothesis on the groups involved, the Genestier–Lafforgue correspondence and the Fargues–Scholze correspondence are compatible with the linkage principle. However, the explicit computation of Tate cohomology groups, in particular, admissibility and the non-vanishing of Tate cohomology $\widehat{H}^{i}(\Gamma,\Pi)$, genericity, are not well understood. We answer some of these questions in this article. In their paper [HV25, Section 5.0.3], Henniart and Vignéras verified Treumann and Venkatesh conjectures for $G(F)={\rm GL}_{2}(F)$ and $G^{\sigma}(F)={\rm SL}_{2}(F)$ with $\ell=2$.
@@
The following are our principal results. The first result of this article is to show that the Tate cohomology groups are finitely generated representations. We prove the following result.
Theorem 1.0.1.
@@
Let $G$ be a connected reductive algebraic group over a non-Archimedean local field $F$. Let $\Gamma\subset{\rm Aut}(G)(F)$ be a subgroup of automorphisms of order $\ell$. Let $(\Pi,V)$ be a finite length $\ell$-modular representation of $G(F)\rtimes\Gamma$. Then the Tate cohomology $\widehat{H}^{i}(\Gamma,\Pi)$ is a finitely generated representation of $G^{\sigma}(F)$.
The above theorem is discussed in Section 2. The proof of Theorem 1.0.1 uses the $\ell$-modular version of Schneider and Stuhler resolutions ([SS97, Theorem II.3.1]) of smooth irreducible representations as in Vignéras’s work [Vig97, Proposition 2.6]. Note that we do not assume anything about $\Gamma$ for the above theorem. However, for the next part on the finite length of the Tate cohomology groups, we restrict to the Galois automorphism case. In what follows, we use the hypothesis and notations of the above theorem.
To prove the admissibility of Tate cohomology groups, we prove that the Jacquet modules of $\widehat{H}^{i}(\Gamma,\Pi)$ are controlled by the Tate cohomology of Jacquet modules of $\Pi$ with respect to $\sigma$-stable parabolic subgroups. This is discussed in Section 3. Let $P$ be a $\sigma$-stable parabolic subgroup of $G$. Then the connected component of the subgroup of $\sigma$-fixed points, denoted by $P^{\sigma}$, is a parabolic subgroup of $G^{\sigma}$. Let $U$ be the unipotent radical of $P$ and let $N$ be the unipotent radical of $P^{\sigma}$. Let $\Psi$ be a character of $U(F)$. We will show that the natural map of twisted Jacquet modules
$\widehat{H}^{i}(\Gamma,\Pi)_{N(F),\Psi}\rightarrow\widehat{H}^{i}(\Gamma,\Pi_{U(F),\Psi}),$ (1.1)
@@
is an injection. This article, in essence, is about various applications of the above injective map. First of which is the following theorem.
Theorem 1.0.2.
@@
Let $G$ be a connected reductive group defined over a non-Archimedean local field $F$, and let $E/F$ be a cyclic Galois extension of degree $\ell$. Let $\Pi$ be a finite length $\ell$-modular representation of $G(E)\rtimes{\rm Gal}(E/F)$. Then the Tate cohomology group $\widehat{H}^{i}({\rm Gal}(E/F),\Pi)$ is a finite length $\ell$-modular representation of $G(F)$.
@@
We prove the above theorem using induction on the dimension of $G$. In the induction argument, we need to control the action of the centre of $G^{\sigma}$ along with the exponents in the Jacquet modules with respect to the parabolic subgroups of $G^{\sigma}$. For this reason we restrict to the case where $\sigma$ is induced by a Galois automorphism. In the Galois automorphism case, for a Levi subgroup $M$ defined over $F$, the action of the centre of $M(F)$ on the ${\rm Gal}(E/F)$-Tate cohomology of the corresponding Jacquet modules of $\Pi$ is via the restriction of the action of the centre of $M(E)$. In other words, we are using the fact that $Z(M(E))\cap G(F)=Z(M(F))$. In the general case, i.e., when $\sigma$ is not necessarily induced by a Galois automorphism, we note that $Z(G^{\sigma})$ can be quite large compared to $Z(G)^{\sigma}$. For instance, $\sigma$ could be an inner automorphism induced by a regular element of order $\ell$, and in this case $G^{\sigma}$ is a torus.
@@
For the next result, proved in Section 4, we assume that $G$ is ${\rm Res}_{E/F}{\rm GL}_{n}$, and $\sigma$ is an automorphism of $G$ induced by a generator of the Galois group ${\rm Gal}(E/F)$. Let $\mathcal{K}$ be the maximal unramified extension of $\mathbb{Q}_{\ell}$ and $\Lambda$ be the ring of integers of $\mathcal{K}$. Let $\Pi$ be an absolutely irreducible, integral representation of ${\rm GL}_{n}(E)$, defined over $\mathcal{K}$ such that $\Pi^{\sigma}\simeq\Pi$. Let $\pi$ be an irreducible $\ell$-adic integral cuspidal representation of ${\rm GL}_{n}(F)$ such that $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lifting of $\pi$. Note that $\Pi$ is not necessarily a cuspidal representation. We show the following result on any $\Lambda[{\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)]$ stable lattice in $\Pi$.
Theorem 1.0.3.
@@
With the Hypothesis on $\Pi$ and $\pi$ as above, let $\mathcal{L}$ be a $\Lambda[{\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)]$ stable lattice in $\Pi$. Then, the Tate cohomology $\widehat{H}^{0}({\rm Gal}(E/F),\mathcal{L})$ is an $\ell$-modular irreducible cuspidal representation of ${\rm GL}_{n}(F)$ and $\widehat{H}^{1}({\rm Gal}(E/F),\mathcal{L})$ is trivial.
@@
In the proof of the above result, we first prove that the ${\rm Gal}(E/F)$-Tate cohomology of a $\Lambda[{\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)]$ stable lattice in $\Pi$ has a unique generic sub-quotient. In fact, this genericity result holds in greater generality. Let $G$ be a split connected reductive group defined over $F$. Let $\Pi$ be a smooth irreducible $\ell$-modular representation of $G(E)$ such that $\Pi$ is stable under the action of ${\rm Gal}(E/F)$. Let $B$ be a Borel subgroup of $G$ defined over $F$, and let $U$ be the unipotent radical of $B$. Let $\psi:U(E)\rightarrow\overline{\mathbb{F}}_{\ell}^{\times}$ be a non-degenerate, ${\rm Gal}(E/F)$ invariant character such that $\psi$ restricts to a non-degenerate character on $U(F)$. If $\Pi$ is $(U(E),\psi)$-generic, then we show that $\widehat{H}^{0}({\rm Gal}(E/F),\Pi)$ has a unique $(U(F),\psi)$-generic sub-quotient (Theorem 4.0.1). Then to prove Theorem 1.0.3, we use the injection (1.1) and prove the right-hand side of the equation (1.1) is trivial by using the fact that the non-trivial Jacquet modules of a non-cuspidal base change lifting $\Pi$ are multiplicity free and their Jordan Holder factors are not stable under ${\rm Gal}(E/F)$.
@@
In Section 5, we then study the Tate cohomology of invariant lattices in representations obtained as Shintani liftings (to ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$) of representations of finite group ${\rm GL}_{n}(\mathbb{F}_{q})$. The analogue of Theorem 1.0.3 is valid for $\Pi$ which is obtained as a Shintani lift to ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ of a cuspidal representation of ${\rm GL}_{n}(\mathbb{F}_{q})$; we use the Deligne–Lusztig character formula to show that, for any $\Lambda[{\rm GL}_{n}(\mathbb{F}_{q^{\ell}})\rtimes{\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})]$ stable lattice $\mathcal{L}$ in $\Pi$, the space $\widehat{H}^{0}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})$ is isomorphic to the Frobenius twist of the $-\ell$ reduction of $\pi$ (Theorem 5.4.3).
@@
We end this article by considering the example of $\theta_{10}$ in Section 7, the unipotent cuspidal irreducible representation of the symplectic group ${\rm Sp}_{4}(\mathbb{F}_{q})$. Let $\pi_{n}$ be the unique cuspidal unipotent representation of ${\rm Sp}_{4}(\mathbb{F}_{q^{n}})$. Note that the $-\ell$ reduction of $\pi_{n}$ is irreducible and it is denoted by $r_{\ell}(\pi_{n})$. For a vector space $W$ over the field $\overline{\mathbb{F}}_{\ell}$, we denote by $W^{(\ell)}$ the $\overline{\mathbb{F}}_{\ell}$-vector space where the scalar action is twisted by the Frobenius automorphism of the field $\overline{\mathbb{F}}_{\ell}$. We have
Theorem 1.0.4.
@@
The above theorem seems to suggest that for all classical groups $G$ admitting a cuspidal unipotent representation, the zeroth Tate cohomology group for the action of ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$ on the $-\ell$ reduction of a cuspidal unipotent representation of $G(\mathbb{F}_{q^{\ell}})$ is isomorphic to the Frobenius twist of the $-\ell$ reduction of the cuspidal unipotent representation of $G(\mathbb{F}_{q})$.
@@
    We use the standard notations attached to any non-Archimedean local field $F$: the ring of integers of $F$ is $\mathfrak{o}_{F}$, the maximal ideal of $\mathfrak{o}_{F}$ is $\mathfrak{p}_{F}$, and the residue field $\mathfrak{o}_{F}/\mathfrak{p}_{F}$ is $k_{F}$. The field $k_{F}$ has $q_{F}$ elements and $q_{F}$ is a power of a prime $p$.
@@
    Let $G$ be a connected reductive algebraic group defined over $F$. Let $\sigma\in{\rm Aut}(G)$, defined over $F$, be an element of prime order $\ell$ and $\ell\neq p$. Let $\Gamma$ be the group generated by $\sigma$. Let $G^{\sigma}$ be the connected component of the fixed points of $\sigma$, and it is assumed to be reductive.
@@
    A representation $(\pi,V)$ of $G(F)$ is said to be smooth if every vector $v\in V$ is fixed by a compact open subgroup of $G(F)$. Let $P$ be an $F$-rational parabolic subgroup of $G$ and the unipotent radical of $P$ is denoted by $U$. Let $M$ be a Levi subgroup of $P$ defined over $F$. For a smooth representation $(\tau,W)$ of $M(F)$, we denote by $i_{P(F)}^{G(F)}(\tau)$ the normalised parabolic induction of $\tau$ to $G(F)$. We use the notation $r_{U(F)}(\pi)$ for the normalised Jacquet module of a smooth representation $(\pi,V)$ of $G(F)$.
@@
    In this article, a smooth representation $(\pi,V)$ of $G(F)$ is called cuspidal if $r_{U(F)}(\pi)=0$, where $U$ is the unipotent radical of any proper $F$-rational parabolic subgroup $P$ of $G$ For a closed subgroup $H_{1}$ of a reductive $p$-adic group $H$, we will denote by ${\rm Ind}_{H_{1}}^{H}$ and ${\rm ind}_{H_{1}}^{H}$ the smooth induction functor and the compact induction functor, respectively.